Dinitrogentetraoxide partially decomposes according to the following equilibrium: N2O4(g) ⇌ 2NO2(g). A 1.00-L flask is charged with 0.04000 mol of N2O4. At equilibrium at 373 K, 0.0055 mol of N2O4 remains. Determine the Keq for this reaction.


1
Expert's answer
2022-04-05T08:46:04-0400

N2O4(g) ⇌ 2NO2(g)

Set up an ICE chart

Initial:

N2O4 = 0.400 M

NO2 = 0 M

Change:

NO2 = +2x

N2O4 = -x

Equilibrium:

NO2 = 2x

N2O4 = 0.400 - x = 0.0055


[N2O4] = x = 0.3945 M

[NO2] = 2x = 2 × 0.3945 = 0.789 M

Keq = [NO2]2 / [N2O4] = (0.789)2 / 0.3945 = 1.578